OpenStudy (anonymous):

A drawer contains 3 tank tops, 6 short-sleeve shirts, 4 three-quarter sleeve shirts, and 5 long-sleeve shirts. What is the probability of randomly choosing a tank top and then a three-quarter sleeve length, if you replace the tank top? My Options are: 7/18 1/27 2/3 1/18

OpenStudy (kropot72):

There is a total of 18 garments in the drawer. The probability of choosing a tank top is 3/18. The probability of choosing a three-quarter sleeve shirt is 4/18. The result of each choice is independent of the other, therefore the required probability is the product of the two values of probability. Required probability is \[\frac{3}{18}\times\frac{4}{18}=\frac{3\times4}{18\times18}=you\ can\ calculate\] Hint: Remember to reduce the result to its lowest terms.

OpenStudy (anonymous):

Thanks so much! This helped a lot! :)

OpenStudy (kropot72):

You're welcome :)